OpenStudy (anonymous):

find an equation for the inverse of the function g(x)=-2x-7

OpenStudy (anonymous):

y=-2x-7 x=-2y-7 x+7=-2y y=(x+7)/(-2) \[g ^{-1}(x)=-x/2 -3.5\]